OpenStudy (anonymous):

5 MEDALS HELP PLEASE A heated piece of metal cools according to the function c(x) = (.5)x - 10, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = -x - 1. What will be the temperature of the metal after five hours?

OpenStudy (anonymous):

Try c(x) + h(x) evaluated at x=5 but wording is odd.

OpenStudy (anonymous):

Select one: a. -6° Celsius b. 26° Celsius ** i think its b c. 32° Celsius d. 38° Celsius

OpenStudy (anonymous):

Odd. h(5) = -5-1=-6 c(5) = 2.5-10 = -7.5 combined = -13.5 Not told how hot the metal was to begin with?